help with solving questions 2 and 3

Solve. 52,3, t 4 =
Answer #1

Given tha 다the co-efreient moksin D 35 20-5)-366 1+806+6) 12+20115 143 -1 | | :2(4-5)-3c2-3)-6(5+6) 3618+56 3-2 | --10+20 + 2(3) ans

Given that

5x^2+3x+4=0

now we find the value of X

x=-3+[3^2-4*5*4]^1/2/2*5

=-3+(9-80)^1/2/10

=-3+8.43/10

=0.543